# Geometric Algebra Meets Large Language Models: Instruction-Based Transformations of Separate Meshes in 3D, Interactive and Controllable Scenes

### v1

@@ Abstract.
journal: TOG
@@
Figure 1. Illustrative demonstration of shenlong, our text-based AI assistant for precise object repositioning in scene edits. Displayed are two edit prompts: on top, moving the tool table to the right; and on the bottom, placing the orange bottle on top of the tool table. The left side of each panel shows the scene before the edit, and the right side shows the scene after the edit. Shenlong processes user text queries and generates the corresponding transformations in terms of Conformal Geometric Algebra (CGA), by harnessing a Large Language Model (LLM). The resulting output is then applied to the scene. For each scene, the user’s prompt and the LLM’s CGA-formatted response are presented. Using CGA instead of baseline representations such as transformation matrices, yields better results as shown in Section 5.2.
1. Introduction
@@
Our primary objective is to enable the instruction-based repositioning of objects within 3D, interactive and controllable scenes. Such rearrangement tasks are defined through various descriptors such as object poses, visual representations, descriptive language, or immersive interactions by agents within the desired end states. In this work, we focus primarily on textual instructions. This activity forms a part of the broader challenge in Embodied AI known as the rearrangement of environments, where the goal is to configure spaces into predetermined states (Batra et al., 2020). The ability to reposition objects accurately and intuitively is essential, especially given its significant implications across fields such as digital entertainment, virtual reality (VR), simulation training, and architectural visualization.
@@
Traditionally, this task has demanded intensive manual effort and specialized knowledge, restricting the efficiency and accessibility of 3D scene editing. Recent developments in Foundational Models (Bommasani et al., 2021) represent a paradigm shift, suggesting that complex scene editing could become more intuitive and accessible through simple text-based instructions. However, existing machine learning-based scene representation techniques such as Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) and Gaussian Splatting (Kerbl et al., 2023) present substantial challenges for precise object repositioning due to their holistic nature, which often obscures individual object details and limits model generalization due to the plurality of possible scenes. (Yu et al., 2020). In contrast, using separate mesh scene representations provides a viable solution by granting direct access to individual mesh components, which facilitates intuitive interactions and precise alignments tailored to specific editing requirements (Zhai et al., 2024). However, the overarching question remains: How effectively can machine learning interact with these separate mesh components for editing?
@@
The advent of Large Language Models (LLMs) (Ramesh et al., 2021; Touvron et al., 2023; Jiang et al., 2023) offers substantial advancements in human-computer interaction by utilizing linguistic proxies to facilitate instruction-based applications (Hong et al., 2023b; De La Torre et al., 2024; Gong et al., 2024; Durante et al., 2024; Yang et al., 2023). Moreover, LLMs have significantly contributed to the field of Neurosymbolic AI (Sheth et al., 2023; Garcez and Lamb, 2023), enhancing our capability to transform linguistic instructions into precise symbolic representations that bridge the gap between LLMs’ comprehensive understanding abilities and the precision required for complex spatial transformations (Dziri et al., 2023; Jignasu et al., 2023; Hong et al., 2023a). This LLM - Neurosymbolic AI integration prompts a critical inquiry into the optimal symbolic notation that encapsulates geometric transformations in a manner that LLMs can readily interpret. Geometric Algebra (GA), also known as Clifford Algebra, provides a robust mathematical structure ideal for managing transformations and interactions of geometric objects, which has been widely applied in Computer Graphics (Papagiannakis, 2013; Papaefthymiou et al., 2016; Gunn and De Keninck, 2019; Hildenbrand and Rockwood, 2022).
@@
In this paper, we use a specific GA model called 3D Conformal Geometric Algebra (CGA) to effectively integrate intuitive linguistic instructions with precise geometric operations, offering a more accessible method for editing separate mesh 3D scenes. This method utilizes the zero-shot learning capabilities of Large Language Models (LLMs), which allows for adaptability to new 3D environments without the requirement for scene-specific training. Our principal contribution is the development of shenlong, a system tailored for separate mesh scene editing, particularly designed for object repositioning through textual descriptions. Integrated within the ThreeDWorld (TDW) framework (Gan et al., 2021), shenlong supports the Unity3D Engine and enhances VR-ready 3D scene interaction capabilities. Through rigorously designed experiments, we show that shenlong markedly surpasses existing LLM-based alternatives, including those used in NVIDIA’s Omniverse, in terms of object repositioning within 3D and interactive scenes. A critical aspect of our work is the detailed examination of the limitations currently faced by LLM solutions in object repositioning tasks. Furthermore, we illustrate that LLMs can proficiently generate and manipulate CGA operations with minimal input. Ultimately, our research marks a significant advancement towards democratizing the creation and manipulation of digital environments by simplifying the complex technicalities traditionally involved in 3D scene editing.
2. Related Work
@@
2.1. Agent-Driven Object Rearrangement
@@
In the realm of object rearrangement, understanding reusable abstractions through geometric goal specifications ranges from simple coordinate transformations to complex multi-object scenarios (Batra et al., 2020). Chang et al. (2023) approach rearrangement as an offline goal-conditioned reinforcement learning problem where actions rearrange objects from an initial setup in an input image to conform to a goal image’s criteria. Kapelyukh and Johns (2023) propose learning a cost function through an energy-based model to favor human-like object arrangements. Simeonov et al. (2023) address rearrangement tasks using Neural Descriptor Fields (Simeonov et al., 2022), assigning consistent local coordinate frames to task-relevant object parts, localizing these frames on unseen objects, and aligning them through executed actions. Furthermore, Zhai et al. (2024) integrate scene graphs with diffusion processes for editable generative models, yet require extensive training and access to the complete scene graph, highlighting the efficiency of our localized transformation approach. Diverging from methods that depend on extensive scene-specific training, our approach leverages the generalization capabilities of LLMs to simplify object repositioning tasks. In contrast to Kwon et al. (2024), who limit their LLM applications to Euclidean spaces for predicting end-effector poses, and Mavrogiannis et al. (2023), who necessitate numerous and cumbersome predefined predicates for translating cooking instructions into Linear Temporal Logic, our method expoits CGA and decomposes repositioning tasks into a minimal set of primitive transformations. Unlike the LLMR approach (De La Torre et al., 2024), which uses specialized LLMs to generate C# Unity code for scene creation and editing, our method avoids direct code generation, thereby ensuring precise scene manipulation without the complexities of writing and debugging code.
2.2. Machine Learning Applications of Geometric Algebra
@@
The incorporation of GA into neural computation was initially introduced in (Pearson and Bisset, 1994), with subsequent developments introducing multivector-valued neurons for radial basis function networks (Corrochano et al., 1996), multilayer perceptrons (MLPs) (Buchholz, 2000; Buchholz and Sommer, 2001), and various neural network architectures (Buchholz and Le Bihan, 2008; Buchholz and Sommer, 2008; Buchholz et al., 2007; Bayro-Corrochano, 2001). Currently, GA-based neural networks have been applied in various domains, including signal processing (Buchholz and Le Bihan, 2008), robotics (Bayro-Corrochano et al., 2018), partial differential equation modeling (Brandstetter et al., 2022), fluid dynamics (Ruhe et al., 2023), and particle physics (Ruhe et al., 2024). Furthermore, novel GA-based architectures such as multivector-valued convolutional neural networks (CNNs) (Li et al., 2022; Wang et al., 2021), recurrent neural networks (Kuroe, 2011; Zhu and Sun, 2016), and transformer networks (Liu and Cao, 2022; Brehmer et al., 2024) have been introduced, demonstrating the versatility and efficacy of GA in enhancing neural network capabilities for geometrically oriented tasks. Unlike these approaches, our work does not incorporate geometric algebraic components directly within deep learning architectures. Instead, we utilize geometric algebra as a communicative mediator between the LLM and our object rearrangement application. This method emphasizes the use of geometric algebra primarily as a tool to enhance the interaction and translation of complex geometrical tasks into understandable formats for the LLM, fostering more effective problem solving capabilities in practical applications. Recently, (Wang et al., 2023) fine-tuned ChatGPT with a large curated collection of textual documents on Geometric Algebra, aimed at developing customized learning plans for students in diverse fields. Our work diverges from the approach of fine-tuning LLMs and instead investigates whether LLMs can effectively utilize geometric algebra for object rearrangement tasks with minimal prompting.
3. Conformal Geometric Algebra
@@
In his seminal 1872 Erlangen Programme, Felix Klein introduced the revolutionary idea that geometry is best explained by algebra, asserting that algebraic structures govern geometric shapes (Hawkins, 1984). Building on this, quaternions have resurged since the late 20th century due to their compact and efficient representation of spatial rotations. GA further unifies and extends algebraic systems, including vector algebra, complex numbers, and quaternions, through the concept of multivectors. Unlike traditional approaches, which treat scalars, vectors, and higher-dimensional entities separately, GA provides a cohesive framework to uniformly represent and manipulate these entities.
@@
At the heart of GA is the concept of the geometric product, which generalizes the dot product and the cross product in Euclidean spaces. Additionally, as all products can be defined using only the geometric one, we need only use the latter one along with addition, scalar multiplication, and conjugation to perform any multivector manipulation. In the context of this work we will be employing CGA, a 32-dimensional extension of quaternions and dual-quaternions (Rooney, 2007), where all entities such as vertices, spheres, planes as well as rotations, translations and dilations are uniformly expressed as multivectors (Kamarianakis and Papagiannakis, 2021).
@@
For example, assuming the standard CGA basic elements $\{e_{1},e_{2},e_{3},\\ e_{4},e_{5}\}$, a basis of CGA would consist of all 32 ($2^{5}$) combinations of up to five of them via geometric product, i.e., $\{1,e_{i},e_{i}e_{j},e_{i}e_{j}e_{k},e_{i}e_{j}e_{k}e_{l},\\ e_{1}e_{2}e_{3}e_{4}e_{5}\text{ for }1\leq i<j<k<l\leq 5\}$. For convenience, we define the vectors $e_{o}=0.5(e_{5}-e_{4})$ and $e_{\infty}=e_{4}+e_{5}$, as well as denote products of basic elements using only subscripts, e.g., $e_{ijk}:=e_{i}e_{j}e_{k}$. Using this notation, a sphere $s$ centered at $x=(x_{1},x_{2},x_{3})$ with radius $r$ amounts to the CGA multivector:
(1) $S=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+\frac{1}{2}(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-r^{2})e_{\infty}+e_{o}.$
@@
(3) $T^{-1}=1+0.5(t_{1}e_{1}+t_{2}e_{2}+t_{3}e_{3})e_{\infty},$
As noted in Kamarianakis and Papagiannakis (2021), a rotation that would normally be expressed by the unit quaternion:
(4) $q:=a-d\boldsymbol{i}+c\boldsymbol{j}-b\boldsymbol{k}.$
@@
Figure 2. Shenlong’s overview: It generates a CGA transformation after templating the query and adds a $\delta$ upward translation for a hypothetical Z-axis collision.
@@
To extract $T,R$ and $D$ from a generated CGA product $M=TRD$, we follow the methodology presented in (Kamarianakis et al., 2024). Specifically, we apply $M$ to a sphere $C$, centered at origin, with radius equal to $1$. The multivector $C^{\prime}$ corresponding to the transformed sphere can be evaluated via the sandwich product $C^{\prime}=MCM^{-1}$, and we can therefore extract its center $x$ and its radius $r$. However since $C^{\prime}$ is the image of the unit, origin-centered, sphere $C$ after applying $M=TRD$, i.e., first a dilation by $D$, then a rotation by $R$ and finally a translation by $T$, its radius should be equal to the scaling factor of $D$ and its center should be the translation vector corresponding to $T$. Since $T$ and $D$ are therefore known, along with $M=TRD$, we can evaluate $R:=T^{-1}MD^{-1}$ and therefore the corresponding unit quaternion using (4) and (5).
4. Methodology
@@
It should be noted that shenlong operates at a local level, lacking a comprehensive understanding of other objects in the scene. While this approach enhances processing efficiency, it may result in collisions with other objects. To address this, we observe that the fuzziness of queries, such as move the ‘sofa’ to the right, allows for multiple possible solutions. Upon detecting a collision, shenlong explores perturbed transformations to resolve the issue, e.g., a $\delta$ upward translation on Z-axis, ensuring minimal deviation from the intended initial transformation. As a final remark, the proposed system is integrated within the ThreeDWorld (TDW) framework (Gan et al., 2021), supports Unity3D Engine and fosters VR-ready 3D scenes, augmenting human VR interaction.
4.1. LLM Processing & CGA Formulation
@@
Through careful prompting, we instruct the LLM on the key facts and operations of CGA. The complete prompt template is provided in Figure 5. Below are some key remarks. We explained how coordinate extraction can utilize the inner product |||| operation to precisely isolate spatial coordinates. For example, $(X_{1}^{\text{max}}|e_{2})$ corresponds to the $y$ coordinate of $X_{i}^{\text{max}}$. Additionally, we defined the outer (or wedge) product, which establishes planes for rotational operations, and details the mechanisms for both translation and rotation rotors. These rotors are not only defined individually but are also combined through rotor composition, enabling complex sequential transformations essential for accurate scene manipulation. To effectively guide the LLM in utilizing the rotation and translation rotors, we provided a total of five illustrative examples. These include a single example each of rotation and translation to demonstrate basic movements, one compositional example that integrates both types of transformations, and two additional examples addressing more complex, fuzzy queries, such as move on top of and move next to another. This small diverse set of examples ensures the LLM effectively understands and implements the necessary algebraic operations for object repositioning. Despite the limited number of examples, i.e., only five, we demonstrate in Section 5.2.3 that the LLM can still generalize to complex queries. Finally, the LLM responds by generating a JSON output that includes the specified rotors for composition to be applied to each object.
4.2. Collision Detection Module

### v2

@@ Abstract
Index Terms:
Large Language Models, Conformal Geometric Algebra, 3D Scene Editing, Separate Meshes, Object Repositioning, Controllable Scene Authoring
I Introduction
@@
Figure 1: Illustrative demonstration of Shenlong, our text-based AI assistant for precise object repositioning in scene edits. Displayed are two edit prompts: on top, moving the tool table to the right; and on the bottom, placing the orange bottle on top of the tool table. The left side of each panel shows the scene before the edit, and the right side shows the scene after the edit. Shenlong processes user text queries and generates the corresponding transformations in terms of Conformal Geometric Algebra (CGA), by harnessing a Large Language Model (LLM). The resulting output is then applied to the scene. For each scene, the user’s prompt and the LLM’s CGA-formatted response are presented. Using CGA instead of baseline representations such as transformation matrices, yields better results as shown in Section V-B.
@@
Our primary objective is to enable the instruction-based repositioning of objects within 3D, interactive and controllable scenes. Such rearrangement tasks are defined through various descriptors such as object poses, visual representations, descriptive language, or immersive interactions by agents within the desired end states. In this work, we focus primarily on textual instructions. This activity forms a part of the broader challenge in Embodied AI known as the rearrangement of environments, where the goal is to configure spaces into predetermined states [1]. The ability to reposition objects accurately and intuitively is essential, especially given its significant implications across fields such as digital entertainment, virtual reality (VR), simulation training, and architectural visualization.
@@
Traditionally, this task has demanded intensive manual effort and specialized knowledge, restricting the efficiency and accessibility of 3D scene editing. Recent developments in Foundational Models [4] represent a paradigm shift, suggesting that complex scene editing could become more intuitive and accessible through simple text-based instructions. However, existing machine learning-based scene representation techniques such as Neural Radiance Fields (NeRFs) [37] and Gaussian Splatting [30] present substantial challenges for precise object repositioning due to their holistic nature, which often obscures individual object details and limits model generalization due to the plurality of possible scenes. [52]. In contrast, using separate mesh scene representations provides a viable solution by granting direct access to individual mesh components, which facilitates intuitive interactions and precise alignments tailored to specific editing requirements [53]. However, the overarching question remains: How effectively can machine learning interact with these separate mesh components for editing?
@@
The advent of Large Language Models (LLMs) [41, 48, 25] offers substantial advancements in human-computer interaction by utilizing linguistic proxies to facilitate instruction-based applications [23, 14, 19, 15, 51]. Moreover, LLMs have significantly contributed to the field of Neurosymbolic AI [45, 18], enhancing our capability to transform linguistic instructions into precise symbolic representations that bridge the gap between LLMs’ comprehensive understanding abilities and the precision required for complex spatial transformations [16, 26, 24]. This LLM - Neurosymbolic AI integration prompts a critical inquiry into the optimal symbolic notation that encapsulates geometric transformations in a manner that LLMs can readily interpret. Geometric Algebra (GA), also known as Clifford Algebra, provides a robust mathematical structure ideal for managing transformations and interactions of geometric objects, which has been widely applied in Computer Graphics [39, 38, 20, 22].
@@
In this paper, we use a specific GA model called 3D Conformal Geometric Algebra (CGA) to effectively integrate intuitive linguistic instructions with precise geometric operations, offering a more accessible method for editing separate mesh 3D scenes. This method utilizes the zero-shot learning capabilities of Large Language Models (LLMs), which allows for adaptability to new 3D environments without the requirement for scene-specific training. Our principal contribution is the development of Shenlong, a system tailored for separate mesh scene editing, particularly designed for object repositioning through textual descriptions. Integrated within the ThreeDWorld (TDW) framework [17], Shenlong supports the Unity3D Engine and enhances VR-ready 3D scene interaction capabilities. Through rigorously designed experiments, we show that Shenlong markedly surpasses existing LLM-based alternatives, including those used in NVIDIA’s Omniverse, in terms of object repositioning within 3D and interactive scenes. A critical aspect of our work is the detailed examination of the limitations currently faced by LLM solutions in object repositioning tasks. Furthermore, we illustrate that LLMs can proficiently generate and manipulate CGA operations with minimal input. Ultimately, our research marks a significant advancement towards democratizing the creation and manipulation of digital environments by simplifying the complex technicalities traditionally involved in 3D scene editing.
II Related Work
@@
II-A Agent-Driven Object Rearrangement
@@
In the realm of object rearrangement, understanding reusable abstractions through geometric goal specifications ranges from simple coordinate transformations to complex multi-object scenarios [1]. [12] approach rearrangement as an offline goal-conditioned reinforcement learning problem where actions rearrange objects from an initial setup in an input image to conform to a goal image’s criteria. [29] propose learning a cost function through an energy-based model to favor human-like object arrangements. [46] address rearrangement tasks using Neural Descriptor Fields [47], assigning consistent local coordinate frames to task-relevant object parts, localizing these frames on unseen objects, and aligning them through executed actions. Furthermore, [53] integrate scene graphs with diffusion processes for editable generative models, yet require extensive training and access to the complete scene graph, highlighting the efficiency of our localized transformation approach. Diverging from methods that depend on extensive scene-specific training, our approach leverages the generalization capabilities of LLMs to simplify object repositioning tasks. In contrast to [32], who limit their LLM applications to Euclidean spaces for predicting end-effector poses, and [36], who necessitate numerous and cumbersome predefined predicates for translating cooking instructions into Linear Temporal Logic, our method expoits CGA and decomposes repositioning tasks into a minimal set of primitive transformations. Unlike the LLMR approach [14], which uses specialized LLMs to generate C# Unity code for scene creation and editing, our method avoids direct code generation, thereby ensuring precise scene manipulation without the complexities of writing and debugging code.
II-B Machine Learning Applications of Geometric Algebra
@@
The incorporation of GA into neural computation was initially introduced in [40], with subsequent developments introducing multivector-valued neurons for radial basis function networks [13], multilayer perceptrons (MLPs) [7, 9], and various neural network architectures [8, 10, 11, 3]. Currently, GA-based neural networks have been applied in various domains, including signal processing [8], robotics [2], partial differential equation modeling [5], fluid dynamics [44], and particle physics [43]. Furthermore, novel GA-based architectures such as multivector-valued convolutional neural networks (CNNs) [33, 50], recurrent neural networks [31, 54], and transformer networks [34, 6] have been introduced, demonstrating the versatility and efficacy of GA in enhancing neural network capabilities for geometrically oriented tasks. Unlike these approaches, our work does not incorporate geometric algebraic components directly within deep learning architectures. Instead, we utilize geometric algebra as a communicative mediator between the LLM and our object rearrangement application. This method emphasizes the use of geometric algebra primarily as a tool to enhance the interaction and translation of complex geometrical tasks into understandable formats for the LLM, fostering more effective problem solving capabilities in practical applications. Recently, [49] fine-tuned ChatGPT with a large curated collection of textual documents on Geometric Algebra, aimed at developing customized learning plans for students in diverse fields. Our work diverges from the approach of fine-tuning LLMs and instead investigates whether LLMs can effectively utilize geometric algebra for object rearrangement tasks with minimal prompting.
III Conformal Geometric Algebra
@@
In his seminal 1872 Erlangen Programme, Felix Klein introduced the revolutionary idea that geometry is best explained by algebra, asserting that algebraic structures govern geometric shapes [21]. Building on this, quaternions have resurged since the late 20th century due to their compact and efficient representation of spatial rotations. GA further unifies and extends algebraic systems, including vector algebra, complex numbers, and quaternions, through the concept of multivectors. Unlike traditional approaches, which treat scalars, vectors, and higher-dimensional entities separately, GA provides a cohesive framework to uniformly represent and manipulate these entities.
@@
At the heart of GA is the concept of the geometric product, which generalizes the dot product and the cross product in Euclidean spaces. Additionally, as all products can be defined using only the geometric one, we need only use the latter one along with addition, scalar multiplication, and conjugation to perform any multivector manipulation. In the context of this work we will be employing CGA, a 32-dimensional extension of quaternions and dual-quaternions [42], where all entities such as vertices, spheres, planes as well as rotations, translations and dilations are uniformly expressed as multivectors [28].
@@
For example, assuming the standard CGA basic elements $\{e_{1},e_{2},e_{3},e_{4},e_{5}\}$, a basis of CGA would consist of all 32 ($2^{5}$) combinations of up to five of them via geometric product, i.e., $\{1,e_{i},e_{i}e_{j},e_{i}e_{j}e_{k},e_{i}e_{j}e_{k}e_{l},e_{1}e_{2}e_{3}e_{4}e_{5}\text{ for }1\leq i<j<k<l\leq 5\}$. For convenience, we define the vectors $e_{o}=0.5(e_{5}-e_{4})$ and $e_{\infty}=e_{4}+e_{5}$, as well as denote products of basic elements using only subscripts, e.g., $e_{ijk}:=e_{i}e_{j}e_{k}$. Using this notation, a sphere $s$ centered at $x=(x_{1},x_{2},x_{3})$ with radius $r$ amounts to the CGA multivector:
$S=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+\frac{1}{2}(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-r^{2})e_{\infty}+e_{o}.$ (1)
@@
$T^{-1}=1+0.5(t_{1}e_{1}+t_{2}e_{2}+t_{3}e_{3})e_{\infty},$ (3)
As noted in [28], a rotation that would normally be expressed by the unit quaternion:
$q:=a-d\boldsymbol{i}+c\boldsymbol{j}-b\boldsymbol{k}.$ (4)
@@
Figure 2: Shenlong’s overview: It generates a CGA transformation after templating the query and adds a $\delta$ upward translation for a hypothetical Z-axis collision.
@@
To extract $T,R$ and $D$ from a generated CGA product $M=TRD$, we follow the methodology presented in [27]. Specifically, we apply $M$ to a sphere $C$, centered at origin, with radius equal to 111. The multivector $C^{\prime}$ corresponding to the transformed sphere can be evaluated via the sandwich product $C^{\prime}=MCM^{-1}$, and we can therefore extract its center $x$ and its radius $r$. However since $C^{\prime}$ is the image of the unit, origin-centered, sphere $C$ after applying $M=TRD$, i.e., first a dilation by $D$, then a rotation by $R$ and finally a translation by $T$, its radius should be equal to the scaling factor of $D$ and its center should be the translation vector corresponding to $T$. Since $T$ and $D$ are therefore known, along with $M=TRD$, we can evaluate $R:=T^{-1}MD^{-1}$ and therefore the corresponding unit quaternion using (4) and (5).
IV Methodology
@@
It should be noted that Shenlong operates at a local level, lacking a comprehensive understanding of other objects in the scene. While this approach enhances processing efficiency, it may result in collisions with other objects. To address this, we observe that the fuzziness of queries, such as move the ‘sofa’ to the right, allows for multiple possible solutions. Upon detecting a collision, Shenlong explores perturbed transformations to resolve the issue, e.g., a $\delta$ upward translation on Z-axis, ensuring minimal deviation from the intended initial transformation. As a final remark, the proposed system is integrated within the ThreeDWorld (TDW) framework [17], supports Unity3D Engine and fosters VR-ready 3D scenes, augmenting human VR interaction.
IV-A LLM Processing & CGA Formulation
@@
Through careful prompting, we instruct the LLM on the key facts and operations of CGA. Below are some key remarks. We explained how coordinate extraction can utilize the inner product ||| operation to precisely isolate spatial coordinates. For example, $(X_{1}^{\text{max}}|e_{2})$ corresponds to the $y$ coordinate of $X_{i}^{\text{max}}$. Additionally, we defined the outer (or wedge) product, which establishes planes for rotational operations, and details the mechanisms for both translation and rotation rotors. These rotors are not only defined individually but are also combined through rotor composition, enabling complex sequential transformations essential for accurate scene manipulation. To effectively guide the LLM in utilizing the rotation and translation rotors, we provided a total of five illustrative examples. These include a single example each of rotation and translation to demonstrate basic movements, one compositional example that integrates both types of transformations, and two additional examples addressing more complex, fuzzy queries, such as move on top of and move next to another. This small diverse set of examples ensures the LLM effectively understands and implements the necessary algebraic operations for object repositioning. Despite the limited number of examples, i.e., only five, we demonstrate in Section V-B3 that the LLM can still generalize to complex queries. Finally, the LLM responds by generating a JSON output that includes the specified rotors for composition to be applied to each object.
IV-B Collision Detection Module